- Status Report
- August 10, 2022
ISS On-Orbit Status 24 June 2002
All ISS systems continue to function nominally,
except as noted previously or below. Starting out on
Week 3, the Expedition 5 crew had a very busy day today.
Workday began with CDR Korzun and FE-2 Treschev readying the Progress
7P cargo ship for undocking. At about 2:30am EDT, Korzun
activated the spacecraft and removed the air duct from the open
Progress/SM hatchway. The two cosmonauts then removed the QD
(quick disconnect) screw clamps of the docking mechanism, while the
SM thrusters were disabled at 2:50am in preparation of the subsequent
leak check. The SM/Progress interface was documented with the
U.S. DVCAM digital camcorder and the video recording downlinked to
the ground via Ku-band. The hatches were then closed and the
leak check conducted. Thrusters were reenabled at 4:40am.
Progress undocking is scheduled for tomorrow morning at 4:26am
EDT, with atmospheric reentry at 8:12am. Progress 8P will then
be launched Wednesday morning at 1:36am EDT, with docking on
Saturday, 6/29, at about 2:30am.
FE-1 Peggy Whitson powered up the EXPRESS Rack 2 (ER2) laptop
computer. The ER2 activation itself was initiated via ground
commanding from the POC (Payload Operations Center), and Peggy’s
action completed the power-up.
Whitson then checked out the MedOps cardiac defibrillator, a routine
task that is performed as soon as possible from Expedition start and
every 60 days thereafter.
After the midday break (8:15-9:15am), CDR Valery Korzun and Peggy
Whitson unstowed the CMRS (crew medical restraint system) from the
CHeCS rack for a checkout. They set the board-like device up by
attaching it at the DDCU rack with the head end toward the CHeCS
(crew health care systems) rack. The CMRS allows strapping down
a patient on the board with a harness for medical attention by the
CMO (crew medical officer), who is also provided with restraints
around the device.
Whitson power-cycled the IV-CPDS (intravehicular charged particle
directional spectrometer), required before the ground can command the
instrument remotely (Mondays and Wednesdays).
The flight engineer then worked on the Zeolite crystal growth (ZCG)
experiment. After the ground had powered it up and checked its
software, she deactivated the CPUs (central processor units) and
their primary and backup hard drive connections of the ZCG IZECS
(improved Zeolite electronic control system). While she had
some time set aside for reviewing and familiarization with the ZCG
configuration, the ground powered down the furnace units, after which
Whitson loaded 19 metal autoclaves into the furnace. Finally,
the alignment guides were removed from the ER2 ARIS (active rack
isolation system) to unlock the rack mechanically and allow active
isolation for the ZCG experiment.
The two cosmonauts completed another session of the Russian-Japanese
HDTV (high-definition TV) experiment, which uses video imaging for
medical evaluation of the crewmember’s close-up facial features.
CDR Korzun spent some time in the DC-1 "Pirs" docking
module with maintenance work on its V1 and V2 fans, cleaning their
FE-2 Sergei Treschev installed the Russian GFI-1
"Relaksatsiya" (Relaxation) experiment by mounting the
ultraviolet (UV) camera with spectrometer unit at SM window #9, after
reconfiguring the Russian payload laptop 3 for the experiment.
[Relaksatsiya deals with the study of the chemoluminescent
chemical reactions and atmospheric light phenomena that occur during
high-velocity interaction between the exhaust products from space
vehicles and the atmosphere at orbital altitude and during the entry
of space vehicles into the Earth’s upper atmosphere. Purpose of
the current experiment is to record the second burn of the
Progress-257 thrusters tomorrow morning with a camcorder. After
Progress’ departure, the experiment will be torn down and stowed
Treschev also spent about an hour in the SM taking noise level
readings with the "Shumometer" sound level meter.
[Purpose of the periodic activity is the characterization of the
potential sources of excessive acoustic noise in the SM cabins and
above the treadmill (ventilation fans located in the RO work
compartment). Only the continuously running vehicle systems in
operation are recorded, with a step-by-step deactivation of
individual ventilation fans. Results of the sampling will be
used in the development of noise reduction measures and in the
validation of the ground acoustic testing of the main noise sources
on RSC-Energia’s SM simulator in KIS-416.]
Peggy Whitson completed the weekly maintenance of the TVIS
(treadmill with vibration isolation and stabilization), while Sergei
Treschev did the monthly maintenance of the CEVIS (cycle ergometer
with vibration isolation).
Regular daily maintenance of the SOSh environment control and life
support system was performed by Valery Korzun, while Whitson did the
autonomous payload status check and Treschev prepared the IMS delta
All crewmembers performed their daily physical exercise.
MCC-H is continuing to research an alarm anomaly noticed by the crew
yesterday with the SM cabin fan caution.
Today’s target areas for the CEO program were Lake Nasser, Toshka
Lakes; Egypt (most of this target area’s features were well to
the left of track for this pass. Of interest: documenting the
water color of the Nile River and Lake Nasser), Lakes of the
Eastern Sierra (although the pass only clipped the lower
corner of this target area, the crew had the opportunity to document
the extremely variable land use patterns from the Los Angeles Basin
to just north of Las Vegas), Medicine Bow NF Fire
(Dynamic Event Site: This 8,600+ acre fire in eastern Wyoming
continues to rage out of control. Of interest: using the long lenses
of the ESC [electronic still camera] to obtain near nadir views of
this event. Oblique, context views, before or after, overpass will
also be helpful), Canadian Rocky Mountains (crew to use
today’s fair weather pass over this target area to obtain context
photos of the region, including oblique views, if possible.
They were also to note the extent of the remaining snow
CEO images can be viewed at the website http://eol.jsc.nasa.gov
U.S. and Russian Segment Status (as of 1:35 pm
Environmental Control and Life Support (ECLSS) and
Thermal Control (TCS):
Elektron O2 generator is powered On (24-amp mode). Vozdukh
CO2 scrubber is ON in MANUAL cycle mode #5 (vacuum pump failed).
U.S. CDRA CO2 scrubber is Off (failed). BMP Harmful
Impurities unit: Absorbent bed #1 in Purify mode, bed #2 in Purify
SM Working Compartment: Pressure (mmHg) — 761,
temperature (deg C) — 27.2, ppO2 (mmHg) — 146.1, ppCO2 (mmHg) —
SM Transfer Compartment: Pressure (mmHg) — 762,
temperature (deg C) — 21.0.
FGB Cabin: Pressure (mmHg) — 756, temperature (deg C)
Node: Pressure (mmHg) — 753.43, temperature (deg C) —
22.2 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
U.S. Lab: Pressure (mmHg) — 755.12, temperature (deg C)
— 25.0, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
Joint Airlock (Equip. Lock): Pressure (mmHg) — 755.02,
temperature (deg C) — 22.5; shell heater temp (deg C) — 23.5, ppO2
(mmHg) — n/a; ppCO2 (mmHg) — n/a.
PMA-1: Shell heater temp (deg C) — 24.0
PMA-2: Shell heater temp (deg C) — 17.0
(Note: Partial pressures ppO2 and ppCO2 in U.S. segment [USOS] not
available because MCA [major constituent analyzer] is failed and in
Extended Life mode [= a state that preserves mass spectrometer
vacuum but produces no pp data]). MSA (mass spectrometer assembly)
and VGA (verification gas assembly) have been removed for return to
Electrical Power Systems (EPS):
Beta Gimbal Assembly (BGA) 2B in Autotrack mode; BGA 4B in
SM batteries: no data available. Battery #1 reported failed
FGB: Battery #2 is offline (ROM mode); battery #1 in
"Cycle" mode. All other batteries (4) are in “Partial
Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby
Thermal Control Systems:
Air conditioner SKV-1 is On. SKV-2 is Off.
Command & Data Handling Systems:
C&C-3 MDM is prime, C&C-2 is back-up, and C&C-1
is in standby.
GNC-1 MDM is prime; GNC-2 is Backup.
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-1 MDM is operational; PL-2 MDM is Off.
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational.
SM Central Computer (TsVM): 2 redundant lanes (of 3) operational.
State vector — US GPS (SIGI string 1)
Attitude — Russian
Angular rates — US RGA1 (rate gyro assembly 1)
All Russian communications & tracking systems are
S-band is operating nominally.
Ku-band is operating nominally.
Audio subsystem operating nominally.
Video subsystem operating nominally.
MCOR (medium-rate communications outage recorder) is powered Off (due
to consistently reaching upper operational temperature during
SSRMS/Canadarm2 at MBS PDGF (mobile base system/power &
data grapple fixture), with Keep Alive power on both strings.
RWS (robotics workstations): Cupola is Active, Lab RWS is Off.
SSRMS Prime string Wrist Roll (WR) joint was replaced on UF-2.
ISS Orbit (as of this morning, 6:51am EDT [=
Mean altitude — 393.1 km
Apogee — 397.8 km
Perigee — 388.4 km
Period — 92.4 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.0006978
Orbits per 24-hr. day — 15.58
Altitude decrease — 130 m (mean) in last 24 hours
Solar Beta Angle — -27.0 deg (magnitude increasing)
Revolutions since FGB/Zarya launch (Nov. ’98) — 20516
Current Flight Attitude — LVLH (local vertical/local horizontal =
“earth-fixed”: z-axis in local vertical, x-axis in velocity vector
[yaw: -10 deg, pitch: -7 deg., roll: 0 deg])
For more on ISS orbit and naked-eye visibility dates/times,